This question involves tobacco use. Men that are non-smokers= 387, Ocassional smokers= 45, Regular Smoker= 90, Heavy Smoker= 37. Women that are non smoker=421, Ocassional smokers=46, Regular smokers =69, Heavy smokers =34.
If one of the 1129 people are randomly selected, find the probability that the person is a man or a heavy smoker. My original answer was .558 and I got it wrong, thank you so much for the help.

Answer by stanbon(75887)

This question involves tobacco use.
Men that are non-smokers= 387, Occasional smokers= 45, Regular Smoker= 90, Heavy Smoker= 37.
Women that are non smoker=421, Occasional smokers=46, Regular smokers =69, Heavy smokers =34.
--------------
If one of the 1129 people are randomly selected, find the probability that the person is a man or a heavy smoker. My original answer was .558 and I got it wrong, thank you so much for the help.
-----
P(man or hs) = P(man) + P(hs) -P(man and hs)
= (559/1129) + (71/1129) - (37/1129) = 0.5252
================================================
Cheers,
Stan H.
===============